### v1

@@ 5 Simulation
    (CLT) $r=3$. $d_{1}=d_{2}=\bar{d}\in\{20,60,100\}$. For each $\bar{d}$, we set $T=200$ and $w_{i}=(r-i+1)\sqrt{d_{1}d_{2}}$. For factor loading vectors, we let $\delta=0.2$ to allow for non-orthogonal loading vectors. The error ${\cal E}_{i,j,t}$ are generated as in Configuration IV to allow for weak cross-sectional correlations. We simulate the distribution of $a_{ik}$ in (26) with $i=1$, $k=1$ under three choices of $u$: $u_{1}=1/\sqrt{\bar{d}}$, $u_{2}=[1,0,0,\ldots,0]^{\top}$ and $u_{3}=[0,1,0,0,\ldots,0]^{\top}$.
Figure 8: QQ plots and histograms of $\sqrt{T}u^{\top}\left(\widehat{a}_{ik}^{{\rm iso}}-\text{sign}(\widehat{a}_{ik}^{{\rm iso}\top}a_{ik})\cdot a_{ik}\right)/\sigma_{u,ik}$ under Configuration VII. Note: 1. The row displays the results for $u_{i}$. The column shows the results for different dimension $\bar{d}$. 2. The red curve plots the distribution of standard normal distribution.
Figure 8 shows the QQ plots and histograms of $\sqrt{T}u^{\top}\left(\widehat{a}_{ik}^{{\rm iso}}-\text{sign}(\widehat{a}_{ik}^{{\rm iso}\top}a_{ik})\cdot a_{ik}\right)/\sigma_{u,ik}$ derived from Theorem 4.3 under Configuration VII. It is observed that the normalized empirical distribution closely approximates the standard normal distribution.
Finally, we evaluate the performance of two proposed rank estimation algorithms: the unfolded eigenvalue ratio method and the eigenvalue ratio method via inner product, over the following DGP configuration:
@@
    (Rank Estimation) $r=3$. $d_{1}=d_{2}=\bar{d}\in\{20,40,60,80\}$ and $T\in\{100,300,500\}$. Set $\delta=0.2$ to allow for correlation among factor loading vectors. Error terms are generated as in Configuration IV to accomodate weak cross-sectional correlation. The factors are generated according to (31) with $w_{i}=(r-i+1)\bar{d}$ and $\phi\in\{0.1,0.5\}$.
@@
6 Empirical Application
6.1 Characteristic decile portfolios
@@
In this section, in line with Babii et al., (2023), we conduct empirical analysis on the dataset collected by Chen and Zimmermann, (2022), consisting of over 200 characteristic-sorted portfolios from previous studies of stock market anomalies. We utilize the August 2023 release of the database, focusing on monthly portfolio returns sorted into 10 deciles based on firm-level characteristics spanning from January 1990 to December 2022. As we only consider a balanced panel of portfolios, the number of characteristics throughout the entire sample period is 127. Therefore, the dimension of the tensor-valued time series ${\cal Y}_{t}$ we considered is $127\times 10$ with sample size $T=396$. Additionally, we obtain the risk-free rate from the Kenneth French data library to compute the excess return of each portfolio.
We rewrite model (1) as:
${\cal Y}_{t,jl}=\sum_{i=1}^{r}f_{it}a_{i1,j}a_{i2,l}+{\cal E}_{t,jl},$ (32)
where ${\cal Y}_{t,jl}$ is the excess return of the $l^{\text{th}}$-decile of the $j^{\text{th}}$ characteristic at time $t=1,\ldots,T$; $f_{it}$ are the systematic risk factors; factor loading $a_{i1,j}$ determines the heterogeneous exposure of the $j^{\text{th}}$ characteristics to the $i^{\text{th}}$ risk factor; loading $a_{i2,l}$ determines the exposure of the $l^{\text{th}}$ decile to the $i^{\text{th}}$ risk factor. In this model, all loading vectors are normalized to $1$, with $f_{it}$ absorbing all scales of loadings. We use the generalized ratio-based method as well as the screen plot to select 3 as the number of factors.
[…]
Table 2: Summary statistics of estimated loadings $\widehat{a}_{i1}$ specific to characteristics
Max Mean Min Std >0
CC-ISO
$\widehat{a}_{1,1}$ 0.065 0.088 0.107 0.009 1
$\widehat{a}_{2,1}$ 0.087 -0.068 -0.212 0.058 0.134
$\widehat{a}_{3,1}$ 0.049 -0.073 -0.188 0.050 0.063
AC-ISO
$\widehat{a}_{1,1}$ 0.038 0.087 0.126 0.019 1
$\widehat{a}_{2,1}$ 0.063 -0.070 -0.208 0.055 0.087
$\widehat{a}_{3,1}$ 0.148 -0.044 -0.230 0.078 0.276
TPCA
$\widehat{a}_{1,1}$ 0.108 0.088 0.064 0.009 1
$\widehat{a}_{2,1}$ 0.275 0.003 -0.230 0.089 0.559
$\widehat{a}_{3,1}$ 0.186 0.006 -0.348 0.089 0.496
GE
$\widehat{a}_{1,1}$ 0.181 0.08 0.001 0.033 1
$\widehat{a}_{2,1}$ 0.245 0.059 -0.087 0.067 0.803
$\widehat{a}_{3,1}$ 0.301 0.006 -0.291 0.089 0.520
Tucker-AC-IP
$\widehat{a}_{1,1}$ 0.14 0.09 0.066 0.01 1
$\widehat{a}_{2,1}$ 0.286 0 -0.295 0.089 0.409
$\widehat{a}_{3,1}$ 0.263 -0.002 -0.187 0.089 0.457
  • Note: $>0$ denotes the ratio of positive entries in each loading vector.
Table 2 reports the summary statistics of the estimated loadings $\widehat{a}_{i1}$, which determine the exposure of characteristics to the three latent factors. The statistics for TPCA closely align with the empirical results documented in Babii et al., (2023). Across all algorithms, the loadings on the first factor are consistently positive with relatively small standard deviations. However, while $\widehat{a}_{2,1}$ and $\widehat{a}_{3,1}$ demonstrate approximate symmetry around zero in TPCA, with approximately half of the loadings being positive, they exhibit significant skewness in the ISO algorithms, with their means deviating from zero. In the case of the GE algorithm, $\widehat{a}_{2,1}$ displays high skewness whereas $\widehat{a}_{3,1}$ shows an approximate symmetry. Specifically, around half of the loadings are positive, and the maximum and minimum values of $\widehat{a}_{3,1,i}$ are approximately symmetric about zero. Similarly, Tucker-AC-IP algorithm also exhibits approximate symmetric patterns in the second factor loading vector.
Table 3: $R$-squared across different algorithms
Method CC-ISO AC-ISO TPCA GE Tucker-AC-IP
$R^{2}$ 0.853 0.844 0.759 0.776 0.878
Table 3 displays the $R^{2}$ of the estimation of Model (32) across five algorithms. In this section, $R^{2}$ is defined as:
@@
where $\tilde{\mathcal{Y}}$ is the demeaned tensor of $\mathcal{Y}$ with the $t^{th}$ entry defined as $\tilde{\mathcal{Y}}_{t}=\mathcal{Y}_{t}-\frac{1}{T}\sum_{t=1}^{T}\mathcal{Y}_{t}.$
Among the four algorithms for the CP factor model, it is observed that the CC-ISO algorithms achieve the best fit to the model, exhibiting the highest $R$-squared values. Comparatively, within the ISO algorithms, CC-ISO outperforms AC-ISO, yielding higher $R$-squared values. However, the TPCA algorithm displays the lowest $R$-squared value at 0.759.
@@
Given that Tucker decomposition entails a larger number of factors compared to CP decomposition, it is not surprising that Tucker-AC-IP yields higher $R^{2}$ than all algorithms based on CP decomposition. Nonetheless, the degree of improvement is modest, with $R^{2}$ increasing by only 2 to 4 percent when adopting the Tucker factor model. This observation suggests that the characteristic decile portfolio data might possess a CP-like factor structure.
Figure 9: $R$-squared of the rolling-sample study on characteristic decile portfolio returns across five algorithms. Note: The first panel shows the results of CC-ISO, TPCA and Tucker-AC-IP. The second panel shows the results of AC-ISO and GE.
We also conduct a rolling-sample study on portfolio excess returns, where each rolling sample spans 120 months, resulting in a total of $T-120=276$ rolling samples with the first rolling sample from January 1990 to December 2000. Within each rolling sample, we estimate the CP factor model in equation (32) using four algorithms and compute the sample $R^{2}$. For comparison, we also calculate the sample $R^{2}$ under tucker tensor factor model using Tucker-AC-IP algorithm .
Figure 9 illustrates the results of the rolling $R^{2}$. The first panel shows the results of the algorithms for the CP factor model based on contemporary covariance matrix, alongside the Tucker-AC-IP algorithm for the Tucker factor model. Meanwhile, the second panel presents the results of algorithms for CP factor model based on auto-covariance matrix.
The three curves in the first panel exhibits a similar pattern: fluctuations with an overall increasing trend between 1990 and 2000, followed by a sharp rise between 2000 and 2003. Subsequently, they revert to fluctuation after 2003. However, the rolling $R^{2}$ of CC-ISO consistently outperforms that of TPCA across all rolling samples. Interestingly, while the Tucker factor model demonstrates the best fit in the full sample, it does not outperform CC-ISO in the rolling sample study, particularly in the rolling samples prior to 2000 and after 2008.
In the second panel, both AC-ISO and GE algorithms show significant fluctuations in $R^{2}$ starting after 2010, which coincides with the COVID pandemic period. Tucker-AC-IP, which also relies on auto covariance matrix estimation, exhibit milder fluctuations during this period. Before 2008, AC-ISO’s $R^{2}$ follows a smoother pattern compared to the GE algorithm.
6.2 Aggregate international trade flow
Understanding the pattern and evolution of international trade flow is essential for a broad range of economic activities including policy-making, economic forecast, and firm-level optimization. In this study, we apply the CP tensor factor model to the international trade flow data, where loading vectors and latent factors are estimated simultaneously.
We use monthly aggregate import and export volumes of commodity goods from January 1991 to December 2015, including a total of $172,800$ observations. The data comes from the International Trade Flow (IMF) Direction of Trade Statistics (DOTS) and involves trade among 24 countries and regions. These include Australia (AU), Canada (CA), China Mainland (CN), Denmark (DK), Finland (FI), France (FR), Germany (DE), Hong Kong (HK), Indonesia (ID), Ireland (IE), Italy (IT), Japan (JP), Korea (KR), Malaysia (MY), Mexico (MX), Netherlands (NL), New Zealand (NZ), Singapore (SG), Spain (ES), Sweden (SE), Taiwan (TW), Thailand (TH), United Kingdom (GB), and the United States (US).
As discussed in Section 2, the latent factors can be interpreted as trading hubs while the factor loadings represent import/export contributions to these hubs. Employing the generalized ratio-based method and the screen plot, we determine 6 as the optimal number of trading hubs.
Table 4 presents the $R^{2}$ across 4 different algorithms for the CP factor model and 2 algorithms for the Tucker factor model: the auto-covariance based algorithm for the Tucker factor model (Tucker-AC-IP) with $h=1$ by Han et al., 2022a and the contemporary-covariance based counterpart (Tucker-CC-IP). In the Tucker factor model, the number of dimensions of the latent hubs is selected as $(4,4)$, following Chen and Chen, (2022). Among the algorithms for the CP factor model, CC-ISO achieves the highest $R^{2}$. AC-ISO follows closely behind, with a slightly lower $R^{2}$ compared to CC-ISO. However, the TPCA algorithm fails to adequately fit the model to the data, resulting in a negative $R^{2}$ value of -2.56.
Table 4: $R$-squared of fitting aggregate international trade flow data across different algorithms
Method CC-PCA AC-PCA TPCA GE Tucker-AC-IP Tucker-CC-IP
$R^{2}$ 0.7951 0.7921 -2.56 0.5782 0.521 0.521
To gain deeper insights into dynamic patterns, we conduct a five-year rolling study on the trade flow data. Each rolling sample spans five years, starting from 1991 through 1995 for the first sample, and progressing consecutively. Within each rolling sample, we estimate the factor loadings using the CC-ISO algorithm proposed in this paper. We assume that the factor loadings remain constant in each rolling sample and fix the number of factors at $r=6$ across all samples. Each sample is indexed by the mid-year of the five-year span.
Unlike the matrix factor model considered in Chen and Chen, (2022), the factor loading vectors in CP factor model are uniquely identified up to the sign change. Therefore, rotation analysis, such as varimax rotation, is not applicable. Instead of applying varimax and interpreting latent hubs by the dominant country/region, we analyze latent hubs based directly on their estimates from the model. In our proposed algorithm, latent hubs are ranked by the corresponding eigenvalues of the unfolded covariance matrix and we fix, with the first latent hub contributing the most to the export/import volumes and variances. However, country contributions to each latent hub can vary across different time periods.
Figure 10 illustrates the relationships between countries and latent hubs for three selected years. The sizes of latent hub nodes are proportional to the strengths of the corresponding factors. The relationships between countries and latent hubs, shown as dotted lines, are plotted using the loading matrix on the export/import side after a truncation transformation. This is achieved by normalizing the loading matrix so that the sum of all entries equals one. Therefore, each entry represents the relative contribution of a country to a latent hub. The figure includes countries with the top four contributions to any latent hubs. Some countries significantly contribute to more than one latent hub, resulting in a stable number of countries on the export/import side, approximately 10. Countries are ranked by their total export/import volume to/from in-sample countries in the selected sub-sample.
As shown in the plot, in 1995, the US had the highest export and import volume, dominating Hub 1 on the export side and Hubs 1 to 4 on the import side. Hub 3 was significantly dominated by China on the export side, with other participants including Japan, Germany, and Taiwan on the export side, and the US, Hong Kong, and Korea on the import side, indicating deep trade connections among these countries. Hub 6 was mainly utilized by European countries on both the export and import sides, reflecting the impact of the foundation of the European Union in 1993.
In 2003, the US remained the country with the highest export and import volumes. However, China surpassed Japan and Germany, becoming the country with the second highest export and import volumes. On the export side, China actively participated in international trade within Hubs 1 to 4, along with the US, Germany, Japan, and European countries. The hub mainly used by European countries experienced growth from 1995 and became the fifth largest hub in 2003. Hub 6 can be interpreted as the APEC hub, as it was primarily used by the US and Asian countries/regions, including Japan, China, Korea, and Taiwan.
By 2012, China had surpassed the US and became the country with the highest trade volume. On the export side, Hubs 1 and 2 were dominated by China. Participants of Hubs 1 and 2 on the import side were mainly from North America and Europe, indicating the growing importance of China in international trade in the 2010s. Additionally, the composition of the hubs became less geographically concentrated: while Hub 6 remained primarily used by European countries on the export side, Canada and Mexico became significant participants in this hub on the import side. Moreover, there was no hub predominantly dominated by Asian countries as in 2003. Japan, Korea, and Singapore were important members of Hub 3, which was shared with the US, Germany, the Netherlands, and France. This suggests that international trade became more global and less regional by 2012.
Figure 10: Trading volume network plots of 3 selected sub-samples. Notes: 1. Relations of the export (import) side between countries and latent hubs are represented on the left (right) half of the figure. 2. On the left(right)-hand side are the union of countries with top 4 contributions to each hub on the export (import) side. 3. Countries are ranked by the total volume of export (left) and import (right) to/from countries in the sample. 4. Sizes of latent hub nodes represent the strengths of the latent hub. 5. Thickness of dotted lines between countries and latent hubs represent the level of contribution.

### v2

@@ 5 Simulation
    (CLT) $r=3$. $d_{1}=d_{2}=\bar{d}\in\{20,60,100\}$. For each $\bar{d}$, we set $T=200$ and $w_{i}=(r-i+1)\sqrt{d_{1}d_{2}}$. For factor loading vectors, we let $\eta=0.1$ to allow for non-orthogonal loading vectors. The error ${\cal E}_{i,j,t}$ are generated as in Configuration IV to allow for weak cross-sectional correlations. We simulate the distribution of $a_{ik}$ in (27) with $i=1$, $k=1$ under three choices of $u$: $u_{1}=1/\sqrt{\bar{d}}$, $u_{2}=[1,0,0,\ldots,0]^{\top}$ and $u_{3}=[0,1,0,0,\ldots,0]^{\top}$.
Figure 8: QQ plots and histograms of $\sqrt{T\Theta_{ii}}u^{\top}\left(\widehat{a}_{ik}^{{\rm iso}}-\text{sign}(\widehat{a}_{ik}^{{\rm iso}\top}a_{ik})\cdot a_{ik}\right)/\sigma_{u,ik}$ under Configuration VII. Notes: (1) The row displays the results for $u_{i}$. The column shows the results for different dimension $\bar{d}$. (2) The red curve plots the distribution of standard normal distribution.
Figure 8 shows the QQ plots and histograms of $\sqrt{T\Theta_{ii}}u^{\top}\left(\widehat{a}_{ik}^{{\rm iso}}-\text{sign}(\widehat{a}_{ik}^{{\rm iso}\top}a_{ik})\cdot a_{ik}\right)/\sigma_{u,ik}$ derived from Theorem 4.4 under Configuration VII. It is observed that the normalized empirical distribution closely approximates the standard normal distribution. Figure 22 in Appendix C shows the corresponding histograms when the variance $\sigma_{u,ik}^{2}$ is replaced by its estimate $\widehat{\sigma}_{u,ik}^{2}$, using the thresholding estimator proposed in Chen et al., (2025). The approximation to the standard normal distribution remains accurate even with estimated variance, demonstrating the robustness of the inference.
Next, we evaluate the performance of two proposed rank estimation algorithms: the unfolded eigenvalue ratio method and the eigenvalue ratio method via inner product, over the following DGP configuration:
@@
    (Rank Estimation) $r=3$. $d_{1}=d_{2}=\bar{d}\in\{20,40,60,80\}$ and $T\in\{100,300,500\}$. Set $\eta=0.1$ to allow for correlation among factor loading vectors. Error terms are generated as in Configuration IV to accommodate weak cross-sectional correlation. Factors are generated according to (32) with $w_{i}=(r-i+1)\bar{d}$ and $\phi\in\{0.1,0.5\}$.
@@
Finally we evaluate the factor estimation of the proposed algorithm and compare it with AC-ISO.
  • IX
    (Factor estimation) Set $r=3$ and $\eta=0.1$. The error ${\cal E}_{i,j,t}$ is generated as in Configuration IV. We vary $(d_{1},d_{2})$ from the set $\{(40,40),(40,60),(60,60)\}$ and $T$ from $\{100,300,500\}$. The estimation error is evaluated by the mean square error:
    $MSE=\frac{1}{rT}\sum_{t=1}^{T}\sum_{i=1}^{r}\left(\widehat{f}_{it}-h_{i}f_{it}\right)^{2},$
    where $h_{i}=w_{i}/\widehat{w}_{i}.$
Figure 9: Boxplots of Factor Estimation Errors under Configuration IX.
Figure 9 shows the results under Configuration IX. The $C$-ISO method outperforms AC-ISO in terms of both the accuracy and variability of the factor estimates. In addition, the estimation errors decrease as $d$ and $T$ increase, which is consistent with Theorem 4.3.
6 Empirical Application: Characteristic decile portfolios
@@
Latent factor models have been extensively employed in the asset pricing literature to identify risk factors and measure risk exposures. In this section, we conduct an empirical analysis using the dataset compiled by Chen and Zimmermann, (2022), which consists of over 200 characteristic-sorted portfolios drawn from previous studies on stock market anomalies. This data has also been analyzed by Babii et al., (2022). We utilize the August 2023 release of the database, which contains monthly portfolio excess returns sorted into 10 deciles based on firm-level characteristics, spanning from January 1990 to December 2022. To ensure a balanced panel of portfolios throughout the sample period, we restrict our analysis to 127 characteristics [^13] The issue of missing values in tensor factor models is both important and challenging. To the best of our knowledge, the only existing study that addresses this issue is Cen and Lam, (2025), which adopts a Tucker tensor factor structure; no prior work has considered missing values under the CP tensor factor framework. Tackling this problem entails substantial additional effort, which we plan to undertake in a separate project.. Consequently, the dimension of the tensor-valued time series ${\cal Y}_{t}$ is $127\times 10$ with a sample size of $T=396$. The risk-free rate, obtained from the Kenneth French data library, is used to compute the excess return of each portfolio.
As noted by Babii et al., (2022), Lettau and Pelger, (2020) also study a similar dataset, which includes 10 single-sorted decile portfolios for each of 37 characteristics. However, rather than utilizing the natural tensor structure, they flatten the data into a total of 370 portfolios and apply risk-premium PCA based on a vector factor model. In contrast, we exploit the inherent tensor structure by estimating a CP tenor factor model via the proposed CC-ISO method. The CP tensor factor model effectively identifies risk factors and factor exposures without rotation ambiguity. This key property ensures that the estimated factors and loadings are uniquely identified up to sign, independent of rotations or the choice of a varimax algorithm, therefore enhancing the reliability of their economic interpretation.
To implement this, we estimate the CP tensor factor model introduced in Equation (1). As discussed in Section 2, the latent factors can be interpreted as systematic risk factors, while the factor loadings represent heterogeneous risk exposures.
To determine the number of factors, we examine the eigenvalues of the sample covariance matrix. The first factor is substantially stronger than the others, with its corresponding eigenvalue accounting for 86% of the total sum of eigenvalues, which might suggest the presence of a dominant market factor. To ensure that our model captures effects beyond this dominant factor, we set the minimum number of factors to two. Guided by the eigenvalue ratio plot and the scree plot (Figure 10), we select three factors for our analysis. For the Tucker factor models, the mode-wise eigenvalue ratio and scree plots (Figure 25 in Appendix D) suggest the rank pair $(2,3)$.
Figure 10: The plot of aggregate eigenvalue ratio (upper panel) and the aggregate scree plot (lower panel) of the excess return of characteristic decile portfolios, starting from the second largest eigenvalue. In the scree plot, the y-axis indicates the proportion of the $i^{th}$ largest eigenvalue to the sum of eigenvalues. In the eigenvalue ratio plot, the y-axis indicates the ratio of the $i^{th}$ eigenvalue to the $(i+1)^{th}$ eigenvalue. The plots suggest 3 factors for our analysis.
We estimate the CP factor model in Equation (1) using various algorithms: CC-ISO, TPCA, AC-ISO with $h=1$, GE proposed by Chang et al., (2023) with $K=1$, the CC-based iterative projection methods based on Tucker decomposition (Tucker-CC-IP) proposed by Han et al., 2022a with rank $(2,3)$, the HOOI algorithm based on Tucker decomposition applied by Lettau, (2024) (3D-PCA), and the observed factor model with Fama-French 3 factors (FF3). We compare the in-sample performance by $R^{2}$ defined as:
@@
where $\widehat{\mathcal{Y}}$ is the fitted value of $\mathcal{Y}$ with various methods and $\tilde{\mathcal{Y}}$ is the demeaned tensor of $\mathcal{Y}$ with the $t^{th}$ entry defined as $\tilde{\mathcal{Y}}_{t}=\mathcal{Y}_{t}-\frac{1}{T}\sum_{t=1}^{T}\mathcal{Y}_{t}.$
To evaluate out-of-sample performance, we conduct one-step-ahead forecasts over the final two years of the sample, following the scheme proposed by Chang et al., (2023). Specifically, for each $s\in\{1,2,\ldots,24\}$, we estimate the model using the sample $\{{\cal Y}_{t}\}_{t=1}^{371+s}$ and fit the estimated latent factors with a VAR(1) model. The one-period-ahead excess return ${\cal Y}_{372+s}$ is then predicted using the estimated loadings and the predicted one-period-ahead factors. The out-of-sample MSE is defined as:
$\text{Out-of-sample MSE}=\frac{1}{24}\sum_{s=1}^{24}\left\|{\cal Y}_{372+s}-\widehat{\cal Y}_{372+s}\right\|_{F}^{2}.$
Table 2 displays the $R^{2}$ and out-of-sample MSE ratios (normalized by Tucker-CC-IP) for six algorithms along with $p$-values from the one-sided Diebold and Mariano, (1995) (DM) Forecast comparison tests. In the DM tests, row “DM I” evaluates whether each competing algorithm outperforms Tucker-AC-IP, while row “DM II” tests whether CC-ISO outperforms the respective competing method. Among the CP-based methods, CC-ISO achieves the best in-sample fit, with the highest $R^{2}$. Within the ISO methods, CC-ISO outperforms AC-ISO. In contrast, The GE algorithm yields the lowest $R^{2}$ value of 0.768.
@@
Given that Tucker decomposition generally involves a much larger number of factors than CP decomposition ($6$ vs $3$ in our example), it is not surprising that Tucker-CC-IP achieves slightly higher $R^{2}$ than the CP-based methods. Nonetheless, the improvement is modest, with Tucker-CC-IP increasing $R^{2}$ by only $2$ to $3$ percentage points compared to CP models. This observation suggests that the characteristic decile portfolio data might exhibit a CP-like factor structure, where a low-rank CP model effectively captures the underlying risk factors.
In terms of forecast performance, smaller MSE ratios indicate better prediction accuracy. CC-ISO significantly outperforms Tucker-CC-IP, 3D-PCA, and GE, while performing comparably to AC-ISO and TPCA. Interestingly, although TPCA achieves the lowest MSE, its improvement over Tucker-CC-IP is not statistically significant. Overall, the Tucker-based methods tend to produce higher MSEs, likely due to overfitting from the increased number of factors. These findings highlight the advantage of CP-based methods in balancing model parsimony with predictive performance.
$C$-ISO AC-ISO TPCA GE Tucker-$C$-IP 3D-PCA FF3
$R^{2}$ 0.8531 0.8445 0.8108 0.7682 0.8783 0.8783 0.839
OOS MSE Ratio 0.9401 0.9471 0.9359 1.0084 1.0000 1.0002 0.9748
DM I 0.0153 0.0738 0.1733 0.6221 - 0.6353 0.1403
DM II - 0.3362 0.5235 0.0083 0.0153 0.0151 0.1165
Table 2: Comparison of different algorithms for characteristic decile portfolio excess returns.
Notes: (1) $R^{2}$: the in-sample $R^{2}$ across all algorithms. OOS MSE ratio: the out-of-sample mean square error ratios relative to Tucker-CC-IP, which has the largest MSE. (2) DM I: one-sided p-value from Diebold and Mariano, (1995) test, with the alternative that the algorithm in each column is more accurate than Tucker-CC-IP. DM II: one-sided p-value from Diebold and Mariano, (1995) test based on the pairwise comparison between the CC-ISO and the algorithm on the column.
Nest, we evaluate cross-sectional pricing errors in the spirit of Lettau, (2024). Specifically, for each decile portfolio, we estimate
$r_{ijt}=\alpha_{ij}+\beta_{ij}^{\prime}f_{t}+u_{ijt},\quad i=1,\ldots,127,j=1,\ldots,10,$
and compute the mean squared pricing errors (MSPE) of $f_{t}$ as
$MSPE_{f}=\frac{1}{127\times 10}\sum_{i,j}\alpha_{ij}^{2}.$
For the Tucker factor model with factor matrix $F_{t}\in\mathbb{R}^{r_{1}\times r_{2}}$ from Tucker factor model, we vectorize it prior to regression, i.e. $f_{t}=\vec{(}F_{t})\in\mathbb{R}^{r_{1}r_{2}}$.
Table 3 reports MSPE and MSPE ratios of different methods relative to Tucker-CC-IP with rank (2,3), as suggested by the mode-wise scree plots. For completeness, we include the 3D-PCA approach of Lettau, (2024) and the Fama-French three-factor model. As shown, the factors obtained from CC-ISO and GE achieve the lowest MSPE, demonstrating that our model performs strongly not only in out-of-sample forecasting but also in cross-sectional pricing.
Method $C$-ISO AC-ISO TPCA GE Tucker-$C$-IP 3D-PCA FF3
MSPE 0.0603 0.0925 0.0632 0.0594 0.0799 0.0785 0.0883
MSPE Ratio 0.7547 1.1573 0.7913 0.7434 1.0000 0.9825 0.8744
Table 3: Table of mean pricing error ratios for different methods. The MSPE row shows the mean squared pricing errors of each method. The MSPE Ratio row shows the mean pricing error ratios relative to Tucker-CC-IP with rank (2,3).
Finally, we further examine the estimated factors and loadings in our CP factor model. The estimated loadings reveal several notable patterns. Figure 11 presents the estimated loadings along with their 95% confidence intervals across characteristics. The loadings on the first factor are uniformly positive, resembling “long-only” portfolio weights. The loadings on the second and third factors display more variation, with predominantly positive elements. These loadings are closely linked to the volatilities of the decile portfolios. In particular, Figure 12 compares the volatilities of decile 10 portfolios with the corresponding characteristic loadings on the third factor. The figure highlights a strong correlation (0.875) between the two: characteristics with higher decile 10 portfolio volatility tend to have larger third-factor loadings, while those with lower volatility have smaller or even negative loadings.
The factor loadings along the decile mode follow a distinct level-slope-curvature pattern, as shown in Figure 13. The first factor follows a “long-only” structure, with strictly positive and nearly uniform weights across deciles. The second factor shows a slope pattern, with decile-10 and decile-1 loadings having opposite signs but similar magnitudes. This structure resembles a “long-short factor”, where high-return deciles are offset by low-return deciles. The third loading shows a convex pattern, forming the “curvature” component.
The consistent “long-only” pattern of the first loading across both modes, combined with the dominant eigenvalue of the first factor, suggests a strong association with the market factor. Figure 14 presents the time series of the first factor along with the excess market return from the Kenneth French data library. The two series exhibit a high correlation (above 0.9), reinforcing the interpretation that this factor primarily captures systematic market risk. In summary, the three estimated factors from the CP factor model can be interpreted as the market factor, the long-short factor, and the volatility factor, respectively.
Figure 11: Estimated loadings on characteristics with 95% confidence interval.
Figure 12: Characteristics loading 3 vs. Decile 10 portfolio volatility.
Figure 13: Estimated loadings on decile with 95% confidence interval.
Figure 14: Factor 1 vs. Excess market return. Factor 1 is scaled so that its volatility is equal to the volatility of the excess market return. The correlation between Factor 1 and the market factor is 0.904.